# StPR: Spatiotemporal Preservation and Routing for Exemplar-Free Video Class-Incremental Learning

### v1

@@ Abstract
Video Class-Incremental Learning (VCIL) seeks to develop models that continuously learn new action categories over time without forgetting previously acquired knowledge. Unlike traditional Class-Incremental Learning (CIL), VCIL introduces the added complexity of spatiotemporal structures, making it particularly challenging to mitigate catastrophic forgetting while effectively capturing both frame-shared semantics and temporal dynamics. Existing approaches either rely on exemplar rehearsal, raising concerns over memory and privacy, or adapt static image-based methods that neglect temporal modeling. To address these limitations, we propose Spatiotemporal Preservation and Routing (StPR), a unified and exemplar-free VCIL framework that explicitly disentangles and preserves spatiotemporal information. First, we introduce Frame-Shared Semantics Distillation (FSSD), which identifies semantically stable and meaningful channels by jointly considering semantic sensitivity and classification contribution. These important semantic channels are selectively regularized to maintain prior knowledge while allowing for adaptation. Second, we design a Temporal Decomposition-based Mixture-of-Experts (TD-MoE), which dynamically routes task-specific experts based on their temporal dynamics, enabling inference without task ID or stored exemplars. Together, StPR effectively leverages spatial semantics and temporal dynamics, achieving a unified, exemplar-free VCIL framework. Extensive experiments on UCF101, HMDB51, and Kinetics400 show that our method outperforms existing baselines while offering improved interpretability and efficiency in VCIL. Code is available in the supplementary materials.
@@ 1 Introduction
Class-Incremental Learning (CIL) li2017learning ; belouadah2021comprehensive ; de2021continual ; masana2022class ; zhang2024center develops models that learn from a sequence of tasks without forgetting previous knowledge, recognizing an ever-growing set of classes without past task data or identifiers. A key challenge is catastrophic forgetting CatastrophicInterference1989McCloskey ; ConnectionistModels1990Ratcliff , where new knowledge overwrites old. While well studied for images, extending CIL to videos—Video Class-Incremental Learning (VCIL) park2021class ; villa2022vclimb —remains underexplored. VCIL differs from CIL by requiring continual learning of new categories while modeling frame-shared semantics and temporal dependencies, unlike CIL’s focus on static images. This spatiotemporal complexity is critical for understanding actions, motion, and scene dynamics in real-world applications like surveillance, driver monitoring, robotics, and lifelogging. Further, memory and privacy constraints often prohibit storing past video data, demanding continual learning without rehearsal.
@@
The central challenge of VCIL lies in mitigating catastrophic forgetting while effectively leveraging frame-shared semantics and temporal dynamics to incrementally learn new categories. Existing methods can be broadly categorized into two types, as illustrated in Figure 1(a): 1) Exemplar-based methods rebuffi2017icarl ; hou2019learning ; douillard2020podnet ; park2021class ; pei2022learning ; villa2022vclimb ; alssum2023just ; liang2024hypercorrelation store a portion of previous data (video clips, frames, or features) and apply rehearsal to reduce forgetting. However, storing exemplars incurs memory and privacy costs and typically emphasizes frame-level learning without explicitly modeling temporal dynamics. 2) CIL-based methods li2017learning ; dhar2019learning ; cheng2024stsp adapt algorithms developed for static images, using techniques like regularization or subspace projection. While avoiding exemplar storage, they often overlook spatiotemporal properties by flattening or underutilizing temporal features. In contrast, our method StPR (Figure 1(b)) explicitly decouples video features into Frame shared semantics and temporal dynamics, and reuses these decomposed components to enhance the model’s ability to adapt continually, thereby reducing forgetting without storing extensive exemplars.
@@
Our framework uniquely bridges the gap between video-specific spatiotemporal representation and class-incremental adaptation. By disentangling and leveraging both spatial semantic channel consistency and temporal dynamics, it offers an effective and explainable solution for continual video understanding. Our main contributions are: 1) We propose a Frame-Shared Semantics Distillation method (FSSD) that preserves frame-shared, semantically aligned spatial channels through semantic importance-aware regularization, optimizing the stability-plasticity trade-off in continual learningl; 2) We design a Temporal Decomposition based Mixture-of-Experts strategy (TD-MoE) that decomposes spatiotemporal features and uses temporal dynamics for expert combination, enabling task-id-free and dynamic adaptation; 3) We present a unified, exemplar-free VCIL framework that achieves state-of-the-art results on UCF101, HMDB51, and Kinetics400, demonstrating the effectiveness of integrating spatial semantics and temporal dynamics in VCIL.
@@ 2.1 Class-Incremental Learning
Class-Incremental Learning (CIL) aims to enable models to continually learn new classes without forgetting previously learned ones. Existing approaches typically fall into three categories: (1) regularization-based methods kirkpatrick2017overcoming ; zenke2017continual ; xiang2022coarse ; zhou2023hierarchical , which constrain parameter updates to preserve prior knowledge, often via knowledge distillation li2017learning ; hou2019learning ; (2) replay-based methods bang2021rainbow ; chaudhry2018riemannian ; rebuffi2017icarl , which store or generate past data to reduce forgetting; and (3) structure-based methods serra2018overcoming ; mallya2018packnet ; mallya2018piggyback ; liang2024inflora ; yu2024boosting , which expand model capacity or isolate task-specific components. Recently, CIL combined with pre-trained vision transformers (ViTs) ermis2022memory ; CODAPromptCOntinual2023Smith ; wang2022dualprompt ; wang2022learning has achieved impressive results by leveraging transferable representations and modularity. Some methods fully fine-tune pre-trained backbones boschini2022transfer ; zhang2023slca , but this is computationally expensive. To address efficiency, parameter-efficient fine-tuning (PEFT) methods have been introduced. Prompt pool-based approaches wang2022learning ; CODAPromptCOntinual2023Smith ; wang2024hierarchical ; zhang2023slca maintain task-specific prompts, while adapter-based methods zhou2023revisiting ; tan2024semantically ; gao2024beyond ; liang2024inflora ; ExpandableSubspace2024Zhou adapt ViTs to new classes with minimal updates. While effective, most CIL strategies were originally developed for static image domains and do not generalize well to video-based scenarios, where temporal dynamics play a critical role.
@@ 2.2 Video Class-Incremental Learning
Action recognition has been widely explored with 2D CNNs using temporal aggregation lin2019tsm ; wang2016temporal and 3D CNNs for joint spatiotemporal modeling carreira2017quo ; tran2015learning . More recent work focuses on improving temporal sensitivity and efficiency feichtenhofer2020x3d ; fan2020rubiksnet . However, these models are trained in static setups and do not address continual adaptation or forgetting.
To address these challenges, Video Class-Incremental Learning (VCIL) extends conventional Class-Incremental Learning (CIL) to spatiotemporal data, introducing additional challenges such as managing temporal variations across tasks. Several recent methods, including TCD park2021class , FrameMaker pei2022learning , and HCE liang2024hypercorrelation , address this setting by storing videos or compressed exemplars. However, these strategies raise concerns related to memory efficiency and data privacy. While SMILE alssum2023just effectively extracts image features from individual frames, it does not explicitly capture temporal information, which may limit its ability to leverage the distinctive decision cues present in video data. Exemplar-free methods such as STSP cheng2024stsp mitigate forgetting via orthogonal subspace projections, but they largely adapt image-domain strategies to video tasks without fully exploiting spatiotemporal structures. In contrast, our approach explicitly decouples and models the spatiotemporal structure of videos. We propose a unified VCIL framework that (i) preserves spatial semantic consistency through Frame-Shared Semantics Distillation (FSSD), enhancing knowledge retention without requiring exemplars, and (ii) leverages temporal dynamics for expert routing via Temporal Decomposition-based Mixture-of-Experts (TD-MoE).
@@ 3 Method
Problem Definition: In the Video Class-Incremental Learning (VCIL) setting, a model is trained across $B$ stages with sequentially arriving datasets $\{\mathcal{D}^{1},\dots,\mathcal{D}^{B}\}$. Each dataset $\mathcal{D}^{b}=\{(V_{j}^{b},y_{j}^{b})\}_{j=1}^{|\mathcal{D}^{b}|}$ corresponds to the $b$-th task, where $V_{j}^{b}$ is the $j$-th video and $y_{j}^{b}$ is its class label. Here, videos primarily represent human action recognition scenarios, where the spatiotemporal dynamics capture motion patterns, subject interactions, and scene context. $|\mathcal{D}^{b}|$ represents the number of samples in the $b$-th task. Let $\mathcal{Y}^{b}$ be the label space of the $b$-th dataset. For all $b\neq b^{\prime}$, the label spaces are disjoint: $\mathcal{Y}^{t}\cap\mathcal{Y}^{b^{\prime}}=\varnothing$. The objective of VCIL is to incrementally train a model over $B$ tasks while maintaining high performance across all accumulated classes $\{\mathcal{Y}^{1},\mathcal{Y}^{2},\dots,\mathcal{Y}^{B}\}$.
@@ 3.1 Spatial and Spatiotemporal Encoder
Spatial Encoder.
The shared adapter module chen2022adaptformer $\mathcal{A}^{b}=\{\mathcal{A}_{l}^{b}\}_{l=1}^{N}$ is utilized with a frozen CLIP-ViT model with $N$ layers of transformer module, serving as the spatial extractor. An adapter is an encoder-decoder architecture embedded into the residual of each transformer layer, facilitating transfer learning and enhancing downstream task performance. Typically, it consists of a down-sampling MLP $\mathbf{W}_{down}\in\mathbb{R}^{d\times d_{h}}$, a ReLU activation $\bm{\phi}(\cdot)$, and an up-sampling MLP $\mathbf{W}_{up}\in\mathbb{R}^{d_{h}\times d}$, where $d$ is the input and output dimension, and $d_{h}$ is the hidden dimension. For adapter, if input is $\mathbf{v}_{i}^{\prime}\in\mathbb{R}^{d}$, which is the $i$-th frame-level feature after the Multi-Head Self-Attention and residual connection in CLIP ViT, the output of adapter is:
@@
More detaiails for Spatial and Spatiotemporal Encoder are provided in appendix B.2 and B.3.
@@ 3.2 Frame-Shared Semantics Distillation
In VCIL, shared adapter modules inevitably drift in feature space when adapting to new tasks, leading to forgetting. Directly applying classic uniform-weighted distillation from CIL to video tasks ignores the differences in semantic importance and temporal stability across video features. This leads to a suboptimal balance between stability and plasticity. To address this, we propose Frame-Shared Semantics Distillation (FSSD), which leverages the unique property of videos: despite variations across frames, core semantics (e.g., action concepts) are consistently captured by a stable subset of channels across frames. By identifying and selectively regularizing these frame-shared informative channels, FSSD preserves prior knowledge while enabling flexible adaptation to new tasks.
@@
where $|\mathcal{C}_{b}|$ is the number of classes in the $b$-th task, and $N_{c}$ is the number of samples per class. $I_{b-1,c,j}$ denotes the frame-shared importance of the $j$-th channel for class $c$ from task $(b{-}1)$. $\bar{V}_{b-1,c,i,j}^{s}$ and $\bar{V}_{b,c,i,j}^{s}$ are the $j$-th channel outputs of the $i$-th sample in class $c$, extracted from the spatial encoders of task $(b{-}1)$ and $b$, respectively, calculated on current data. As the previous model is frozen during training, estimating semantic importance is essential for preserving prior knowledge across tasks.
@@
where $\mu_{c,j}$ and $\sigma_{c,j}^{2}$ denote the mean and variance across frames. The Fisher Information $\mathcal{I}(\mu_{c,j})$ (Detailed derivations are provided in the appendix B.1) with respect to $\mu_{c,j}$ is computed as:
$\mathcal{I}(\mu_{c,j})=\mathbb{E}\left[\left(\frac{\bar{V}^{s}_{c,j}-\mu_{c,j}}{\sigma_{c,j}^{2}}\right)^{2}\right]=\frac{1}{\sigma_{j}^{4}}\mathbb{E}\left[(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}\right]=\frac{1}{\sigma_{j}^{4}}\cdot\sigma_{c,j}^{2}=\frac{1}{\sigma_{c,j}^{2}}.$ (5)
@@
where $\|\bar{\mathbf{V}}^{s}_{c}\|\approx\lambda$ is treated as a constant after normalization.
Combining semantic sensitivity and classification score, the semantic importance for the $j$-th channel of the $c$-th class is defined as:
@@
After each task, we compute the frame-shared semantic importance for all classes and accumulate it as the distillation weight for the next task. FSSD ensures that, when learning new tasks, the model retains channels that are more semantically important for previous tasks while allowing less critical parameters to adapt to new knowledge, achieving a better balance between stability and plasticity.
@@ 3.3 Temporal Decomposition based Mixture-of-Experts
Given the high forgetting tendency of deep transformers and the unknown number of future tasks in VCIL, we allocate a dedicated spatiotemporal encoder for each task. As task ID are unavailable during inference, we introduce a routing mechanism to dynamically assign weights to the experts. This routing leverages the temporal patterns extracted from the video features, ensuring relevant experts contribute more to the final representation. This approach enables task ID-agnostic, exemplar-free inference, effectively leveraging the temporal structure of action videos to alleviate forgetting.
Task-Specific Expert. For each task, we train a dedicated expert based on the spatiotemporal encoder. The spatiotemporal features $\mathbf{V}^{st}\in\mathbb{R}^{d_{vt}}$ captured by each expert are computed as in Eq. 2. During inference, we introduce a dynamic routing mechanism, assigning weights to each expert based on temporal relevance, which allows us to obtain the Mixture-of-Experts (MoE) model feature.
Temporal Decomposition base-Router.
@@
For temporal residuals, we observe that redundant frames, where backgrounds and subjects remain consistent, cause minimal variation between adjacent frames kim2024fusion ; liu2021no . This leads to short-term temporal stationarity, allowing frame features to be decomposed into a shared static component and small temporal residuals. We express each frame feature as $\mathbf{V}_{i}^{s}=\bar{\mathbf{v}}+\bm{\epsilon}_{i}$, where $\bar{\mathbf{v}}$ captures the shared static components (e.g., background, actor) and $\bm{\epsilon}_{i}$ represents zero-mean temporal fluctuations. The spatial representation is then the mean across frames:
@@
For the inter-frame information, since the spatiotemporal feature $\mathbf{V}^{st}$ is computed by the attention module, it can be approximated as $\mathbf{V}^{st}\approx\sum_{i=1}^{N_{f}}a_{i}\cdot\mathbf{V}^{s}_{i}$, where $a_{i}$ is attention score. After normalization, we can obtain $\sum_{i=1}^{N_{f}}a_{i}=1$. Substituting this into the decomposition, we can simplify the expression in Eq. 8 as:
@@
$\mathbf{V}^{tem}=\sum_{j=1}^{N_{f}}\left(a_{i}-\tfrac{1}{N_{f}}\right)\cdot\bm{\epsilon}_{i}.$ (10)
This formulation reveals that $\mathbf{V}^{tem}$ quantifies the deviation between the model’s attention-weighted temporal dynamics and the uniform temporal mean. It isolates temporal variations from static spatial semantics, allowing the model to use temporal cues for routing decisions while avoiding background interference, thus supporting our goal of mitigating forgetting and leveraging temporal dynamics for continual learning.
@@
$\mathbf{V}=\bar{\mathbf{V}}^{s}+\sum_{k}r_{k}\cdot\mathbf{V}_{k}^{cls}.$ (12)

### v2

@@ Abstract
Video Class-Incremental Learning (VCIL) seeks to develop models that continuously learn new action categories over time without forgetting previously acquired knowledge. Unlike traditional Class-Incremental Learning (CIL), VCIL introduces the added complexity of spatiotemporal structures, making it particularly challenging to mitigate catastrophic forgetting while effectively capturing both frame-shared semantics and temporal dynamics. Existing approaches either rely on exemplar rehearsal, raising concerns over memory and privacy, or adapt static image-based methods that neglect temporal modeling. To address these limitations, we propose Spatiotemporal Preservation and Routing (StPR) mechanism, a unified and exemplar-free VCIL framework that explicitly disentangles and preserves spatiotemporal information. We begin by introducing Frame-Shared Semantics Distillation (FSSD), which identifies semantically stable and meaningful channels by jointly considering channel-wise sensitivity and classification contribution. By selectively regularizing these important semantic channels, FSSD preserves prior knowledge while allowing for adaptation. Building on this preserved semantic space, we further design a Temporal Decomposition-based Mixture-of-Experts (TD-MoE), which dynamically routes task-specific experts according to temporal dynamics, thereby enabling inference without task IDs or stored exemplars. Through the synergy of FSSD and TD-MoE, StPR progressively leverages spatial semantics and temporal dynamics, culminating in a unified, exemplar-free VCIL framework. Extensive experiments on UCF101, HMDB51, SSv2 and Kinetics400 show that our method outperforms existing baselines while offering improved interpretability and efficiency in VCIL.
@@ 1 Introduction
Class-Incremental Learning (CIL) li2017learning ; belouadah2021comprehensive ; de2021continual ; masana2022class ; zhang2024center develops models that learn from a sequence of tasks without forgetting previous knowledge, recognizing an ever-growing set of classes without past task data or identifiers. A key challenge is catastrophic forgetting CatastrophicInterference1989McCloskey ; ConnectionistModels1990Ratcliff , where new knowledge overwrites old. While well studied for images, extending CIL to videos: Video Class-Incremental Learning (VCIL) park2021class ; villa2022vclimb , remains underexplored. VCIL differs from CIL by requiring continual learning of new categories while modeling frame-shared semantics and temporal dependencies, unlike CIL’s focus on static images. This spatiotemporal complexity is critical for understanding actions, motion, and scene dynamics in real-world applications like surveillance, driver monitoring, and robotics. Further, memory and privacy constraints often prohibit storing past data, demanding continual learning without rehearsal.
@@
The central challenge of VCIL lies in mitigating catastrophic forgetting while effectively leveraging frame-shared semantics and temporal dynamics to incrementally learn new categories. Existing methods can be broadly categorized into two types, as illustrated in Figure 1(a): 1) Exemplar-based methods rebuffi2017icarl ; hou2019learning ; douillard2020podnet ; park2021class ; pei2022learning ; villa2022vclimb ; alssum2023just ; liang2024hypercorrelation ; chen2025csta store a portion of previous data (video clips, frames, or features) and apply rehearsal to reduce forgetting. However, storing exemplars incurs memory and privacy costs and typically emphasizes frame-level learning without explicitly modeling temporal dynamics. 2) CIL-based methods li2017learning ; dhar2019learning ; cheng2024stsp adapt algorithms developed for static images, using techniques like regularization or subspace projection. While avoiding exemplar storage, they often overlook spatiotemporal properties by flattening or underutilizing temporal features. In contrast, our method StPR (Figure 1(b)) explicitly decouples video features into Frame shared semantics and temporal dynamics, and reuses these decomposed components to enhance the model’s ability to adapt continually, thereby reducing forgetting without storing extensive exemplars.
@@
Our framework uniquely bridges the gap between video-specific spatiotemporal representation and class-incremental adaptation. By disentangling and leveraging both spatial semantic channel consistency and temporal dynamics, it offers an effective and explainable solution for continual video understanding. Our main contributions are: 1) We propose a Frame-Shared Semantics Distillation method (FSSD) that preserves frame-shared, semantically aligned spatial channels through semantic importance-aware regularization, optimizing the stability-plasticity trade-off in continual learningl; 2) We design a Temporal Decomposition based Mixture-of-Experts strategy (TD-MoE) that decomposes spatiotemporal features and uses temporal dynamics for expert combination, enabling task-id-free and dynamic adaptation; 3) We present a unified, exemplar-free VCIL framework that achieves state-of-the-art results on UCF101, HMDB51, SSv2 and Kinetics400, demonstrating the effectiveness of integrating spatial semantics and temporal dynamics in VCIL.
@@ 2.1 Class-Incremental Learning
Class-Incremental Learning (CIL) aims to enable models to continually learn new classes without forgetting previously learned ones. Existing approaches typically fall into three categories: (1) regularization-based methods kirkpatrick2017overcoming ; zenke2017continual ; xiang2022coarse ; zhou2023hierarchical , which constrain parameter updates to preserve prior knowledge, often via knowledge distillation li2017learning ; hou2019learning ; (2) exemplar-based methods bang2021rainbow ; chaudhry2018riemannian ; rebuffi2017icarl , which store or generate past data to reduce forgetting; and (3) structure-based methods serra2018overcoming ; mallya2018packnet ; mallya2018piggyback ; liang2024inflora ; yu2024boosting , which expand model capacity or isolate task-specific components. Recently, CIL combined with pre-trained vision transformers (ViTs) ermis2022memory ; CODAPromptCOntinual2023Smith ; wang2022dualprompt ; wang2022learning has achieved impressive results by leveraging transferable representations and modularity. Some methods fully fine-tune pre-trained backbones boschini2022transfer ; zhang2023slca , but this is computationally expensive. To address efficiency, parameter-efficient fine-tuning (PEFT) methods have been introduced. Prompt pool-based approaches wang2022learning ; CODAPromptCOntinual2023Smith ; wang2024hierarchical ; zhang2023slca maintain task-specific prompts, while adapter-based methods zhou2023revisiting ; tan2024semantically ; gao2024beyond ; liang2024inflora ; ExpandableSubspace2024Zhou adapt ViTs to new classes with minimal updates. While effective, most CIL strategies were originally developed for static image domains and do not generalize well to video-based scenarios, where temporal dynamics play a critical role.
@@ 2.2 Video Class-Incremental Learning
Action recognition has been widely explored with 2D CNNs using temporal aggregation lin2019tsm ; wang2016temporal and 3D CNNs for joint spatiotemporal modeling carreira2017quo ; tran2015learning . More recent work focuses on improving temporal sensitivity and efficiency feichtenhofer2020x3d ; fan2020rubiksnet . However, these models are trained in static setups and do not address continual adaptation or forgetting. To address these challenges, Video Class-Incremental Learning (VCIL) extends conventional Class-Incremental Learning (CIL) to spatiotemporal data, introducing additional challenges such as managing temporal variations across tasks. Several recent methods, including TCD park2021class , FrameMaker pei2022learning , and HCE liang2024hypercorrelation , address this setting by storing videos or compressed exemplars. However, these strategies raise concerns related to memory efficiency and data privacy. While SMILE alssum2023just effectively extracts image features from individual frames, it does not explicitly capture temporal information, which may limit its ability to leverage the distinctive decision cues present in video data. Exemplar-free methods such as STSP cheng2024stsp mitigate forgetting via orthogonal subspace projections, but they mainly adapt image-domain strategies to video tasks. In contrast, our approach decouples and models the spatiotemporal structure of videos, proposing a unified VCIL framework that preserves spatial consistency via Frame-Shared Semantics Distillation (FSSD) for knowledge retention without exemplars, while leveraging temporal dynamics for expert routing through Temporal Decomposition-based Mixture-of-Experts (TD-MoE).
@@ 3 Method
Problem Definition: In the Video Class-Incremental Learning (VCIL) setting, a model is trained across $B$ stages with sequentially arriving datasets $\{\mathcal{D}^{1},\dots,\mathcal{D}^{B}\}$. Each dataset $\mathcal{D}^{b}=\{(V_{j}^{b},y_{j}^{b})\}_{j=1}^{|\mathcal{D}^{b}|}$ corresponds to the $b$-th task, where $V_{j}^{b}$ is the $j$-th video and $y_{j}^{b}$ is its class label. Here, videos primarily represent human action recognition scenarios, where the spatiotemporal dynamics capture motion patterns, subject interactions, and scene context. $|\mathcal{D}^{b}|$ represents the number of samples in the $b$-th task. Let $\mathcal{Y}^{b}$ be the label space of the $b$-th dataset. For all $b\neq b^{\prime}$, the label spaces are disjoint: $\mathcal{Y}^{b}\cap\mathcal{Y}^{b^{\prime}}=\varnothing$. The objective of VCIL is to incrementally train a model over $B$ tasks while maintaining high performance across all accumulated classes $\{\mathcal{Y}^{1},\mathcal{Y}^{2},\dots,\mathcal{Y}^{B}\}$.
@@ 3.1 Spatial and Spatiotemporal Encoder
Spatial Encoder. The shared adapter module chen2022adaptformer $\mathcal{A}^{b}=\{\mathcal{A}_{l}^{b}\}_{l=1}^{N}$ is utilized with a frozen CLIP-ViT model with $N$ layers of transformer module, serving as the spatial extractor. An adapter is an encoder-decoder architecture embedded into the residual of each transformer layer, facilitating transfer learning and enhancing downstream task performance. Typically, it consists of a down-sampling MLP $\mathbf{W}_{down}\in\mathbb{R}^{d\times d_{h}}$, a ReLU activation $\bm{\phi}(\cdot)$, and an up-sampling MLP $\mathbf{W}_{up}\in\mathbb{R}^{d_{h}\times d}$, where $d$ is the input and output dimension, and $d_{h}$ is the hidden dimension. For adapter, if input is $\mathbf{v}_{i}^{\prime}\in\mathbb{R}^{d}$, which is the $i$-th frame-level feature after the Multi-Head Self-Attention and residual connection in CLIP ViT, the output of adapter is:
@@ 3.2 Frame-Shared Semantics Distillation
In VCIL, shared adapter modules inevitably drift in feature space when adapting to new tasks, leading to forgetting. Directly applying classic uniform-weighted distillation from CIL to video tasks ignores the differences in semantic importance and temporal stability across video features. This leads to a suboptimal balance between stability and plasticity. To address this, we propose Frame-Shared Semantics Distillation (FSSD), which identifies stable cross-frame channels capturing core semantics and regularizes them to preserve prior knowledge while allowing adaptation
@@
where $|\mathcal{C}_{b}|$ is the number of classes in the $b$-th task, and $N_{c}$ is the number of samples per class. $I_{b-1,c,j}$ denotes the frame-shared importance of the $j$-th channel for class $c$ from task (b1)(b{-}1). $\bar{V}_{b-1,c,i,j}^{s}$ and $\bar{V}_{b,c,i,j}^{s}$ are the $j$-th channel outputs of the $i$-th sample in class $c$, extracted from the spatial encoders of task (b1)(b{-}1) and $b$, respectively, calculated on current data, with the previous model frozen.
@@
where $\mu_{c,j}$ and $\sigma_{c,j}^{2}$ denote the mean and variance across frames. The Fisher Information $\mathcal{I}(\mu_{c,j})$ (Detailed derivations are provided in the appendix B) with respect to $\mu_{c,j}$ is computed as:
$\mathcal{I}(\mu_{c,j})=\mathbb{E}\left[\left(\frac{\bar{V}^{s}_{c,j}-\mu_{c,j}}{\sigma_{c,j}^{2}}\right)^{2}\right]=\frac{1}{\sigma_{c,j}^{4}}\mathbb{E}\left[(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}\right]=\frac{1}{\sigma_{c,j}^{4}}\cdot\sigma_{c,j}^{2}=\frac{1}{\sigma_{c,j}^{2}}.$ (5)
@@
where $\|\bar{\mathbf{V}}^{s}_{c}\|\approx\lambda$ is treated as a constant after normalization. Combining semantic sensitivity and classification score, the semantic importance for the $j$-th channel of the $c$-th class is defined as:
@@
FSSD accumulates frame-shared semantic importance as distillation weights, retaining key channels for old tasks while allowing less important ones to adapt, thus balancing stability and plasticity.
@@ 3.3 Temporal Decomposition based Mixture-of-Experts
Given the high forgetting tendency of deep transformers in VCIL, we allocate a dedicated spatiotemporal encoder for each task. As task IDs are unavailable during inference, we allocate a spatiotemporal encoder per task and design a routing mechanism that dynamically weights experts based on temporal patterns, ensuring relevant experts contribute more to the final representation.
Task-Specific Expert. For each task, we train a dedicated expert based on the spatiotemporal encoder. The spatiotemporal features $\mathbf{V}^{st}\in\mathbb{R}^{d_{vt}}$ captured by each expert are computed as in Eq. 2.
Temporal Decomposition-based Router.
@@
Figure 3: For each video, we uniformly sample 8 frames to compute the $p$-value defining $p>0.05$ as strictly stationary and $p>0.02$ as weakly stationary in the short term.
For temporal residuals, we observe that redundant frames, where backgrounds and subjects remain consistent, cause minimal variation between adjacent frames kim2024fusion ; liu2021no . This leads to short-term temporal stationarity, which we further validate on the UCF101 and HMDB51 datasets Fig. 3. Thus, each frame feature is decomposed as $\mathbf{V}_{i}^{s}=\bar{\mathbf{v}}+\bm{\epsilon}_{i}$, with $\bar{\mathbf{v}}$ as shared static components and $\bm{\epsilon}_{i}$ as temporal residuals. The spatial representation is then the mean across frames:
@@
For the inter-frame information, since the spatiotemporal feature $\mathbf{V}^{st}$ is computed by the attention module, it can be approximated as $\mathbf{V}^{st}\approx\sum_{i=1}^{N_{f}}a_{i}\cdot\mathbf{V}^{s}_{i}$, where $a_{i}$ is attention score. After normalization, we can obtain $\sum_{i=1}^{N_{f}}a_{i}=1$. Substituting Eq. 8, we obtain:
@@
$\mathbf{V}^{tem}=\sum_{i=1}^{N_{f}}\left(a_{i}-\tfrac{1}{N_{f}}\right)\cdot\bm{\epsilon}_{i}.$ (10)
This formulation reveals that $\mathbf{V}^{tem}$ quantifies the deviation between the model’s attention-weighted temporal dynamics and the uniform temporal mean, effectively disentangling temporal variations from static semantics. This enables routing to exploit temporal cues while avoiding background interference, thereby mitigating forgetting and enhancing continual learning.
@@
$\mathbf{V}=\bar{\mathbf{V}}^{s}+\sum_{k}r_{k}\cdot\mathbf{V}_{k}^{st}.$ (12)